# Vulnerabilities that arise from poor governance in Distributed Ledger Technologies

### v1

@@ Abstract.
Current implementations of governance in Distributed Ledger Technologies leave them susceptible to a number of attacks. We survey the state of the art of Distributed Ledger Technologies’ (DLTs) governance protocols and work carried out to systematise good governance properties in the context of DLTs. We then select the most appropriate taxonomy of good governance properties and point to formal security notions that good governance protocols should satisfy. We point practitioners to existing solutions to deliver them, where possible. Furthermore, we outline a number of vulnerabilities that arise in the absence of good governance properties. We call on the research community and DLT research practitioners to prioritise delivering these good governance properties and continue to develop tools to do so, to avoid attacks to DLT protocols that exploit their poor governance models.
@@ 1. Introduction
Distributed Ledger Technologies (DLTs) have enabled a paradigm shift in how users interact and transact online. DLTs are comprised of users, that interact with the DLT network and receive some service from it, and service providers (miners and node operators) that maintain the DLT network by running its protocol and providing service to its users. DLTs do not exist in a single server, but rather run on a web of connected nodes, operated by the service providers, which communicate following a protocol. A protocol is simply a set of instructions that nodes execute to enable the functioning of the DLT. DLTs are run through code, and maintainable code requires updates. DLTs must have a mechanism through which protocols are updated, bugs are patched and the service is improved. This component is crucial to ensure DLTs adapt, develop and improve in a sustainable manner, as well as providing a much needed protection against threats. Without this ability to mutate, vulnerabilities cannot be patched. Since all nodes must agree to use the same protocol, a DLT must rely on some mechanism utilising the underlying consensus algorithm to decide which protocol updates will take place and when. Governance is an all-encompassing umbrella term that captures this ability: the decision-making process to change a DLT protocol.
Governance is (one of) the Achilles’ heel of DLTs. It is central to the promise of decentralisation and it is difficult to untangle. There are many different research lenses from which it can be approached, and often they cannot be divorced from one another. Governance is at the intersection of numerous research areas: legal, software, political, institutional, philosophical, cryptographic, social choice, economic and financial are just some examples. Extensive literature exists outlining parallels with DLT governance and other organisational models such as DAOs, or outlining desirable properties, however, we find it particularly challenging to extract tangible solutions from existing work. Indeed, there is literature outlining what are desireable properties of governance, but little guidance on how to actually achieve and implement said criteria and what are the practical risks when these are not satisfied. We note that there has been a recent upsurge in the research of DAO governance, and we argue that parallels may be drawn from DAO and DLT governance. This follows naturally from the fact that DAOs are often built on DLTs.
Our research focus is on security and its implications in different aspects of the governance protocol. We point practitioners on what existing solutions are available to achieve certain governance properties, and what are the security risks of failing to do so. We outline which desirable properties are currently mutually exclusive, and which need not be.
@@ 2. Related Work
A comprehensive systematization of knowledge on good governance properties can be found in (Kiayias and Lazos, 2022) and (Liu et al., 2023). The authors in (Kiayias and Lazos, 2022) focus on governance properties, whereas (Liu et al., 2023) answer questions on what governance is, who is involved, when is it used and why. The work in (Kiayias and Lazos, 2022) serves as a starting place for the work we present. Namely, we find their taxonomy of governance properties the most clear, tangible and concise thus far. However, they explicitly note that they do not provide specific techniques, algorithms or mechanisms to achieve good governance properties.
@@
In terms of vulnerabilities that may arise as a consequence of poor governance in DLTs, (Schneider, 2022) claims that economic incentives alone do not suffice to enable good governance, and makes a case that well studied social choice and politics findings should be applied to governance in DLTs. The work in (Benedict, 2019) outlines a number of challenges regarding DLT governance. Namely, it notes how poor governance models can cause oligarchies in DLTs, how incentives are crucial in ensuring accountability and notes how decision rights should be clearly defined in the governance protocol. The authors in (De Filippi and Loveluck, 2016) identify that governance in Bitcoin is not as decentralised and code-driven as expected, and uncovers that in fact the decision-making power is highly technocratic and determined by a small minority of code developers. None of the aforementioned works address the formal security and privacy notions necessary in the governance protocol, or the vulnerabilities that arise in the absence of these. We find a substantial overlap in the study of governance vulnerabilities in DAOs and DLTs. The work in (Austgen et al., 2023) and (Daian et al., 2018) outline how vote buying cartels may be automated as a consequence of poor governance models in DAOs. A noteworthy body of research is carried out in (Tan et al., 2023), where the authors note the importance of classical security notions for e-voting should also hold in DAO governance mechanisms. We agree. However, they do not provide guidance on how to achieve these primitives. In summary, some recent works have emerged outlining vulnerabilities that arise as a consequence of poor governance models in DAOs, but less so regarding DLT governance. Some cautionary warnings have been raised regarding DLT governance, but we find no work that explores the suite of possible attacks that may arise in DLTs in the absence of standard good governance properties.
@@ 3. Governance Components
In this work, we refer to governance as the decision-making processes to enable the functioning and updating of a DLT protocol. Before proceeding with necessary criteria that good governance should satisfy, we first outline what are the crucial components of any governance protocol:
@@
  • The vote aggregation algorithm.
  • The election officials. Note that these are participants in charge of tallying the votes, maintaining the voting platform and publishing results.
  • The bulletin board. This is the digital record of cast ballots.
@@
  • The voting mechanism. This describes how votes are tallied to return a winner.
  • The platform to vote.
DLTs allow for novel ways to implement and deliver some of these components. This opens a wide range of opportunities, but like any new and unexplored field, may come with landmines. To understand these risks, we begin with reiterating the taxonomy of governance properties first introduced in (Kiayias and Lazos, 2022). The reader should note the subtle distinction: the following are properties of the aforementioned governance components.

### v2

@@ Abstract.
Distributed Ledger Technologies (DLTs) promise decentralization, transparency, and security, yet the reality often falls short due to fundamental governance flaws. Poorly designed governance frameworks leave these systems vulnerable to coercion, vote-buying, centralization of power, and malicious protocol exploits—threats that undermine the very principles of fairness and equity these technologies seek to uphold. This paper surveys the state of DLT governance, identifies critical vulnerabilities, and highlights the absence of universally accepted best practices for good governance. By bridging insights from cryptography, social choice theory, and e-voting systems, we not only present a comprehensive taxonomy of governance properties essential for safeguarding DLTs but also point to technical solutions that can deliver these properties in practice. This work underscores the urgent need for robust, transparent, and enforceable governance mechanisms. Ensuring good governance is not merely a technical necessity but a societal imperative to protect the public interest, maintain trust, and realize the transformative potential of DLTs for social good.
@@ 1. Introduction
Distributed Ledger Technologies (DLTs) have enabled a paradigm shift in how users interact and transact online. DLTs are comprised of users, that interact with the DLT network and receive some service from it, and service providers (miners and node operators) that maintain the DLT network by running its protocol and providing service to its users (Romero Ugarte, 2018). DLTs do not exist in a single server, but rather run on a web of connected nodes, operated by the service providers, which communicate following a protocol (Sunyaev and Sunyaev, 2020). A protocol is simply a set of instructions that nodes execute to enable the functioning of the DLT. DLTs are run through code, and maintainable code requires updates (Liu et al., 2020). DLTs must have a mechanism through which protocols are updated, bugs are patched and the service is improved. This component is crucial to ensure DLTs adapt, develop and improve in a sustainable manner, as well as providing a much needed protection against threats (Accenture and the Depository Trust & Clearing Corporation , DTCC). Without this ability to mutate, vulnerabilities cannot be patched. Since all nodes must agree to use the same protocol, a DLT must rely on a consensus algorithm to decide which protocol updates will take place and when. Governance is an all-encompassing umbrella term that captures this ability: the decision-making process to change a DLT protocol (Lehdonvirta and Ali, 2016).
Whilst some earlier work laid the foundations (Vishnumurthy et al., 2003), (Dai, 1998), (Chaum, 1983), (Back et al., 2002) it is the Bitcoin white paper (Nakamoto, 2008) that is widely regarded to have kickstarted the Web 3 revolution. Satoshi Nakamoto proposes a peer-to-peer electronic cash system in the wake of the Great Recession of 2008. Blockchain was conceived as a reactionary and idealistic attempt to redistribute power and decentralise decision-making of financial service providers [^1] Satoshi states in 2009: “The root problem with conventional currency is all the trust required to make it work. The central bank must be trusted not to debase the currency, but the history of fiat currencies is full of breaches of that trust. Banks must be trusted to hold our money and transfer it electronically, but they lend it in waves of credit bubbles with barely a fraction in reserve. We have to trust them with our privacy, trust them not to let identity thieves drain our accounts. Their massive overhead costs make micropayments impossible.” (Nakamoto, 2009)..
Cryptocurrencies initially faced a shy uptake but exploded in value after 2017. At time of writing (November 2024), Bitcoin price is up by 24,801.30% since it was created. Similarly, Ether increased by 1,985.73%, Cardano by 929.34% and Solana by 37,300% (Forbes, 2024). Cryptocurrencies are here to stay: governments have not only begun to regulate these assets (examples include the US (Hughes, 2017), Japan (Arora, 2020) and the EU (Zetzsche et al., 2021)), but also own and invest in them as well (Mitchell, 2024) (El Salvador (Gaikwad and Mavale, 2021) and Brazil (Redman, 2024) actively investing, and other countries holding seized assets such as the US and UK are just some examples (Lee, 2024)).
Although Satoshi Nakamoto conceived Bitcoin as a decentralised, peer-to-peer network, the reality is that it and many other cryptocurrencies are centralised (Beikverdi and Song, 2015) (Sai et al., 2021), manipulable (Dhawan and Putniņš, 2023) (Barnes, 2018) and can be undermined due to their governance protocols (Daian et al., 2019), (Judmayer et al., 2019). Governance is critical to delivering the promise of decentralisation, and without good governance, the vulnerabilities that may arise are dire. These issues must be addressed to preserve the original purpose of using cryptocurrencies for social good. Cryptocurrencies have become a matter of public interest: they are used by governments, their combined market worth is of \$3.71 Trillion (Forbes, 2024) and the entry barrier to invest in them is very low, meaning many with low technical and financial literacy can and do participate and invest in these technologies (Bartlett, 2019). Whilst this is revolutionary, it is also highly dangerous: many victims have lost life’s savings due to scams, hacks, heists and rug pulls with little to no compensation (Stanford, 2021). Governance safeguards are necessary to ensure transparency, accountability and prevent monopolisation and theft of funds.
Our work addresses the governance dimension of DLTs that is increasingly intersecting with public policy, regulation, and societal well-being. Cryptocurrencies and distributed ledger technologies (DLTs) are no longer niche financial tools; they are evolving into systems that challenge traditional governance models and require new frameworks for oversight, trust-building, and public protection. By identifying vulnerabilities and proposing governance best practices, this work contributes to the interdisciplinary dialogue on how digital technologies transform governance structures, impact public institutions, and necessitate regulatory adaptations. Ultimately, we hope this research informs policy-making, aids regulators in understanding systemic risks, and helps ensure that these technological innovations serve, rather than exploit, the public interest.
2. Overview
This paper is organized as follows: Section 1 introduces the motivation for examining governance in Distributed Ledger Technologies (DLTs). Section 3 reviews related work, identifying gaps in the current understanding and implementation of governance properties in DLTs and the vulnerabilities they face. Section 4 provides a detailed overview of our methodology, Section 5 introduces the governance components needed to outline our main findings. Subsequently, we describe vulnerabilities that may arise due to current governance methods in DLTs in Section 6, and we present a taxonomy of governance properties in Section 7. The latter also includes possible technical solutions that may be implemented to deliver good governance properties. Section 8 evaluates a selection of DLT applications against our framework of governance properties, and identifies gaps and mutually exclusive properties. Finally, Section 9 concludes with a summary of our contributions and directions for future research.
@@ 3. Related Work
A comprehensive systematization of knowledge on good governance properties can be found in (Kiayias and Lazos, 2022) and (Liu et al., 2023a). The authors in (Kiayias and Lazos, 2022) focus on governance properties, whereas (Liu et al., 2023a) answer questions on what governance is, who is involved, when is it used and why. The work in (Kiayias and Lazos, 2022) serves as a starting place for the work we present. Namely, we find their taxonomy of governance properties the most clear, tangible and concise thus far. However, they explicitly note that they do not provide specific techniques, algorithms or mechanisms to achieve good governance properties.
@@
In terms of vulnerabilities that may arise as a consequence of poor governance in DLTs, (Schneider, 2022) claims that economic incentives alone do not suffice to enable good governance, and makes a case that well studied social choice and politics findings should be applied to governance in DLTs. The work in (Benedict, 2019) outlines a number of challenges regarding DLT governance. Namely, it notes how poor governance models can cause oligarchies in DLTs, how incentives are crucial in ensuring accountability and notes how decision rights should be clearly defined in the governance protocol. The authors in (De Filippi and Loveluck, 2016) identify that governance in Bitcoin is not as decentralised and code-driven as expected, and uncovers that in fact the decision-making power is highly technocratic and determined by a small minority of code developers. None of the aforementioned address the formal security and privacy notions necessary in the governance protocol, or the vulnerabilities that arise in the absence of these. We find a substantial overlap in the study of governance vulnerabilities in Decentralised Autonomous Organisations (DAOs) and DLTs. The work in (Austgen et al., 2023) and (Daian et al., 2018) outline how vote buying cartels may be automated as a consequence of poor governance models in DAOs. The authors in (Feichtinger et al., 2024) also highlight a number of governance related vulnerabilities in DAOs. Specifically regarding purchaseable voting power, malicious proposals and flash loans exploits, which the work in (Kharman and Smyth, 2024) also discusses. A noteworthy body of research is carried out in (Tan et al., 2023), where the authors note the importance of classical security notions for e-voting should also hold in DAO governance mechanisms. We agree. However, they do not provide guidance on how to achieve these primitives. In summary, some recent works have emerged outlining vulnerabilities that arise as a consequence of poor governance models in DAOs, but less so regarding DLT governance. Some cautionary warnings have been raised regarding DLT governance, but we find no work that explores the suite of possible attacks that may arise in DLTs in the absence of standard good governance properties.
4. Methodology
Our work employs a qualitative, analytical approach to identify vulnerabilities in Distributed Ledger Technologies (DLTs) arising from deficiencies in governance models.
4.1. Identification of Vulnerabilities
From work previously carried out regarding DAO vulnerabilities in (Kharman and Smyth, 2024), we identified their governance protocols enable a number of weaknesses as a consequence of flaws in their governance protocols. Other researchers corroborate these findings: (Feichtinger et al., 2024), (Austgen et al., 2023), (Tan et al., 2023). We hypothesize that these vulnerabilities are also applicable to DLTs, since DAOs are implementations built on DLTs. We then proceeded to survey existing work following two criteria: governance related vulnerabilities, and governance properties of DLTs.
4.2. Identification of Governance Properties and Vulnerabilities
A comprehensive review of existing literature and frameworks on governance in DLTs was conducted to identify key governance vulnerabilities and properties. This was done following the PRISMA methodology (Moher et al., 2010). We searched 4 databases using the keywords: DLT, governance, vulnerabilities, attacks, properties. The outcome of this search is summarised in Figure 1.
Figure 1. PRISMA flow diagram for systematic review of literature
The number of works identified was limited, we therefore widened our search to include DAO governance, of which only the work of (Tan et al., 2023), (Feichtinger et al., 2024) and (Daian et al., 2018) were identified as relevant. Of the surveyed studies, only (Kiayias and Lazos, 2022) focuses on governance properties of DLTs, and (Gojka et al., 2021) on vulnerabilities of DLTs, although it does not explicitly discuss which of these vulnerabilities arise as a consequence of poor governance. The authors of (Tan et al., 2023) point to some formal notions of good governance in the context of DAOs, and implementations for governance that attempt to deliver these. These, in conjunction with (Kiayias and Lazos, 2022) serve as the starting point for our work regarding our contributions on governance properties, and (Daian et al., 2018) and (Gojka et al., 2021) for our contributions on governance related vulnerabilities.
4.3. Assessment of Governance Gaps
Using the governance properties as a benchmark, an analysis of governance models in various DLT systems was performed. This step involved examining case studies, technical documentation, and academic studies to identify instances where these governance properties were insufficiently implemented or entirely absent.
4.4. Linking Vulnerabilities to Absent Governance Properties
The identified governance failings were analysed to determine their impact on DLTs. Logical reasoning and case-based evidence were used to connect missing governance properties to specific vulnerabilities, such as vote buying, coercion, centralisation of power, incentives to heist the DLT and non-compliance with election outcomes.
4.5. Theoretical Justification
We then leveraged existing theories and frameworks in organizational governance, distributed systems, e-voting, cryptography and social choice theory to substantiate the connections between governance deficiencies and the emergence of vulnerabilities.
@@ 5. Governance Components
In this work, we refer to governance as the decision-making processes to enable the functioning and updating of a DLT protocol. The authors in (Zachariadis et al., 2019) similarly state that the term governance refers to ‘the frameworks and processes that oversee decision-making, control, and the evolution of the system. This encompasses how rules are established, enforced, and modified within the decentralized network’. The work in (Lehdonvirta and Ali, 2016) state that governance in DLTs is the ‘rule-making by the owners or participants of a system with the purpose of safeguarding their private interests’. Other approaches examine the intersection between blockchain governance and corporate governance (Yermack, 2017) and borrow from the definitions of corporate governance (Davidson, 2024). For example, the work in (Yermack, 2017) states: ‘Ultimately blockchains must rely on a governance process in which the users agree upon a set of requirements for the underlying software code to be changed, including provisions for dispute resolution, sanctions for violating the agreed upon rules, and procedures for enforcement of penalties’. Whilst there is no consensus on what is the definition of governance in the context of DLTs, most definitions overlap with the following critical functionalities: consensus, rules and decision-making.
Before proceeding with the outlined vulnerabilities, we establish definitions and nomenclature to aid with the explanation of governance related attacks. We present what we consider to be the minimum viable components to enable decision-making on-chain. These are borrowed from electronic voting literature, that also establish similar roles (Adida, 2008), (Chaum et al., 2008), (Clarkson et al., 2008), (Ryan et al., 2009).
@@
  • The vote tallying algorithm.
  • The election officials: these are participants in charge of tallying the votes, maintaining the voting platform and publishing results.
  • The bulletin board: the digital record of cast ballots.
@@
  • The platform to vote: the server, website, chain or other online location where votes are cast.
6. Vulnerabilities
6.1. Exploits leveraging public votes
The right to a private vote is recognized as a human right as per Article 21.3 of the Universal Declaration of Human Rights (UN General Assembly, 1948) in 1948. In the absence of it, voters may be coerced, bribed or sell their votes. Secret ballots are congruent to freedom, they ensure voters can express their will free of external influences and pressures. This is no different in DLT governance.
6.1.1. Voter Coercion
When ballots are public, everyone can know how a voter voted. This provides verifiability: everyone can check that the election outcome is correct and that only those allowed to vote voted. However, this forgoes freedom. Free will cannot be protected when a voter faces the reality of not abiding with their coercer’s threats (Lever, 2007). The coercer can easily check if their victim complied. Voters may chose to sell their votes even in the absence of coercion. In some cases, DLTs provide a unique set of conditions where the most beneficial strategy for some voters is to sell their vote. Other works have identified this issue in DAO governance (Daian et al., 2018), (Austgen et al., 2023). We highlight that this problem is prevalent in DLT governance too.
We stress, at current time of writing, none of these solutions offer secret ballots. No existing solution can be shown to satisfy formal notions of ballot secrecy, because the ledger is public. Some have attempted to provide increased levels of privacy, but this privacy is short lived, and often compromised to achieve verifiability. An example is Snapshot’s recently added feature of ‘shielded voting’, wherein votes remain private only until the end of the election. Then, ‘when voting ends, all votes are revealed - as well as who voted for what’  [^2] Snapshot adds ‘Shielded Voting’.. This still enables voter coercion and vote selling.
6.1.2. Disincentive to vote against whales
Both users and miners vote with their feet on a DLT. When a proposal arises for a protocol update, both must agree for it to be implemented. Users may favour a certain proposal, but miners can refuse to service transactions following it. Similarly, if miners favour a proposal that users disagree with, the users may simply stop sending transactions on that chain, and miners will earn no fees. Let us consider a voter with much larger voting power than the rest a whale. An example could be a user with a large share of tokens or a collective with high mining power (a large mining pool) in the DLT. When a whale votes for a certain proposal, other agents are incentivised to follow suit. Otherwise a fork [^3] A fork is a divergence in a DLT regarding the accepted version of transactions (Wang et al., 2019). may arise, and eventually one of the two will outgrow the other and become the main chain. Miners processing transactions on the losing fork will not receive their reward. Consequently users’ transactions will not get approved, and thus they too will eventually leave the fork. Because whales hold greater voting power, the fork they vote for will be the winning one with greater likelihood. Less influential voters will minimise their chance of losses by following the behaviour of whales, which is public, on-chain. This creates a herding effect, and a disincentive to vote true to one’s beliefs, if these oppose those of a whale’s.
6.1.3. Voter Information Asymmetry
Another simpler exploit consists of an attacker purchasing enough voting power to approve a malicious proposal in their favour. Public voting facilitates this attack: the attacker knows exactly how much voting power they need to acquire to overturn an election in their favour. These attacks have been witnessed in DAOs a number of times  [^4] Build DAO’s hostile governance takeover attack, Feb 2022.,  [^5] Beanstalk cryptocurrency project robbed after hacker votes to send themselves \$182 million.,  [^6] Sanctioned Tornado Cash DAO governance heisted by hacker., and are plausible in DLTs too. As shown in (Wang et al., 2020) and (Gencer et al., 2018), miners’ voting power (hash rate distribution) is highly centralized. In some cases, the voting power held by one agent can suffice to fork a network. In others, malicious users may leverage flash loans [^7] Loans written in smart contracts that enable participants to quickly borrow funds without the need for collateral (Chainalysis, 2023). to rapidly obtain sufficient tokens, submit a proposal to their benefit, cash out and repay the loan, whilst still making a profit. Miners are incentivised to service this proposal: they stand to gain high transaction fees. Indeed, the behaviour of whales is known to affect the price of a token (Manahov, 2022): they have the power to send large scale transactions that may incentivize miners to act against the consensus protocol (Liao and Katz, 2017), (Daian et al., 2020a).
6.2. Exploit leveraging non-binding elections
We note that, although informal voting channels may exist, these are in fact nothing other than a polling mechanism, because the only binding votes on a DLT protocol update are the transactions sent by users and approved by miners on-chain, following either one protocol or another. All other off-chain so-called voting platforms are operating under the assumption that miners and users will behave on-chain in the same way that they expressed they would off-chain. The New York Agreement is a clear example of how off-chain voting is not binding and how this assumption may not always hold [^8] Bitcoin’s New York Agreement was a planned upgrade of Bitcoin to increase the block size from 1MB to 2MB. The update garnered significant community support but large Bitcoin mining pools refused to implement it (Harper, 2019). This update would have lowered miner’s transaction fees by improving transaction efficiency. The result was that Bitcoin was forked: Bitcoin maintained the same block size and Bitcoin Cash, the fork, adopted the increased block size (Webb, 2018). (Hunter, 2022). Anyone can vote, but if the agents responsible for implementing the election outcome disagree with the outcome, they have no obligation to implement it, even if the election was fair, free and correct. This is a vulnerability, as stakeholders may subvert an election outcome by refusing to comply. To prevent this, DLT governance protocols must be binding. To ensure this, elections must be carried out on-chain. On-chain voting brings a suite of challenges with it, which we outline in section 6.3.2.
6.3. Exploits leveraging purchasable voting power
6.3.1. Plutocracy
1-dollar-1-vote mechanisms and any variant of these where voting power may be purchased will converge towards granting power to the most wealthy. Any governance mechanism that allows users to purchase voting power will inherently converge to plutocracy and centralise over time (Reijers et al., 2021). Whilst this issue is well known, DLTs continue to implement token based voting, for lack of a better alternative for a decentralised Sybil-protection [^9] Sybil-protection mechanisms are those that prevent an attacker from creating multiple fake identities to gain unfair advantages in a network (Mohaisen and Kim, 2013). system. DLT practitioners must compromise and choose which of the alternatives suits their application needs best: permissionless voting does not rely on a central trusted service to prevent Sybil attacks, and is often implemented by providing voting rights proportional to wealth or computational power. However, it is known to centralise voting power in the hands of the wealthy. Permissioned voting introduces reliance on a centralised trusted authority to issue voting rights, or requires a practical and functional decentralised identity management solution. This permissioned alternative is often implemented through reputation based voting schemes, soulbound tokens, NFTs, or a Know Your Customer solution, checking the validity of participants’ identity.
6.3.2. On-Chain Voting
6.4. Other vulnerabilities
6.4.1. Bribe Automation
This exploit is a consequence of the following properties being absent: ballot secrecy, non-purchaseable voting power and fee-less voting. In DLTs where mining is competitive, if one mining pool gains a transaction, the other do not, so we can model the mining pools to be explicitly distinct from each other, and the smallest mining pool is of size 1 agent. Public votes make it evident to see how a mining pool voted by simply looking up the on-chain transactions. Less influential miners should either align their vote to those of the largest mining pools, or risk having their transactions end up in a fork and not receive the reward for the work they performed to include it. It is therefore easy for a briber to identify miners of less influence and bribe them to ‘vote’ for their desired protocol. The small sized mining pools will make a profit, and the briber can identify how many miners they must bribe to succeed in implementing their desired protocol. As the authors in (Austgen et al., 2023) showed, paradoxically, the more decentralised the network, the easier it is to create a systematic bribing cartel. Small sized miners are cheaper to bribe than large mining pools. The authors even provide a protocol to automate this vote buying in DAOs. This can be re-purposed to buy a miner’s vote in a DLT. The bribe can be paid out as soon as the smart contract detects a transaction following the desired protocol of the briber.
6.4.2. Submitting Malicious Protocol Updates
Malicious protocol updates are not infrequent, and not only in DAOs or DLTs [^10] Backdoor found in Library update, used by major Linux Distributions.. Some go undetected due to low participation rates in governance, in others, the protocol update is forcibly approved by a whale, and in other occasions, the malicious protocol is hidden or not evidently detectable by the community. The consequences of this can be dire: double spending attacks, delaying governance outcomes, creating new tokens or loss of funds are amongst the many attack vectors possible.
It is not feasible to characterise all the ways in which malicious protocol updates may be submitted. Governance is only one of the components necessary in the DLT to prevent these exploits, but a robust governance protocol does not guarantee protection against malicious protocol updates. Access control mechanisms, security and trust assumptions are other components that must be taken into consideration. We encourage DLT design practitioners to make choices in their governance protocol that will alleviate the risk of this exploit. Namely, removing incentives to submit malicious proposals, or to comply with bribes to implement them.